From nch@roe.ac.uk Mon Jan 25 15:36:59 2010 Date: Fri, 22 Jan 2010 17:44:31 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting, 22/01/10 Minutes of WFAU VDFS Science Archive meeting: January 22nd 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, PMW, RPB, MSH, MAR, NJC, RGM, RSC Apologies: JPE, JMS, AL DoNM: 10am, Friday January 29th 2010 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: NCH to remove any RequiredStacks definitions from VSA_InitiateArchive.sql Discharged ACTION: NJC, MAR and ETWS to draft PSPI VSA presentations by Fri Jan 22. Discharged; the team agreed to meet at 3:30pm in the Vista Hut to dry-run/review. Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB (with advice/help from RSC and ETWS) to do a VSA forward-curate test with the ingested V0.6 data. Continues (see below) Actions carried forward from 15/01/10 meeting: ---------------------------------------------- ACTION: RPB to check out availability/maturity of LTO4 drives and then liaise with PMW over purchase if required. Specific points and new actions: -------------------------------- Project management: Interviews for manager/technical-lead post take place on Tuesday 26th; RGM and AL (along with John Peacock and Arthur Trew) are on the panel while PMW and NCH will meet-and-greet the candidates. RGM noted that he needs some updated stats for the submission to STFC concerning VISTA exploitation/added-value; some info was gathered during the meeting and otherwise RGM will be asking various folks for input. WFCAM & VISTA updates: No significant news this week. Comments and issues arising from CASU minutes: None this week. Networking: ETWS noted that all transfers are up-to-date and worked fine for WFCAM and VISTA data. ETWS suggested that the UKLight connection be moved from djedefre to the beefier khafre to test out parallel WFCAM/VISTA transfers. MSH said this should be no problem. WSA/VSA Operations: RPB noted that source merging for the VSA is rather sluggish; turns out that seaming indexes are missing in the schema. RSC is sorting this out. ETWS noted that ingestion of the latest VISTA data was slower than normal for the last batch because of a couple of missing archive keywords in the FITS files... hence only SV Orion has been pushed forward for a catalogue DB release. It is hoped that at least some of the dry-run datasets can be similarly curated to test release DBs early next week in time for the PSPIs meeting. ETWS noted the SW system upgrades to menkaure and khafre are complete, but the team noted some issues with third-party memory leeks (Python?) running under Debian "lenny" so offending modules have had to be reverted to older versions. It turns out one part of the SV Orion survey is synoptic/deep with multiple exposures in the same field. RSC noted that this should be dealt with via the usual ProgrammeBuilder procedure once the deep stack/tile CASU code for VIRCam is in place. Hardware and Systems: ETWS suggested we upgrade the curation servers by procuring a new machine and retire and older, slower model; MSH volunteered to look into specs for a new machine. PMW noted that a few responses to a tender for a new 0.1 PB (yes, really) NAS box had been received, with the usual results at both high and low bids. Software: RSC noted: "Tested our curation code against the latest systems software upgrade and made the necessary updates. Excluded cross-semester non-surveys from being released to allow the release of the 09A non-surveys. Investigated some memory allocation & open file limit issues in the Python software following the systems software upgrade that cropped up during the 09A release." MAR noted that the UI has been fettled to make VSA proprietary protection more WSA-programme like for individual survey protection. Survey Data Release: RSC asked about the status of DR7; it was noted that QC1 awaits eyeball results. Concerning the GPS, it's unknown what the requirements for reprocessing may be. ACTION: NCH to ask GPS SH Phil Lucas about eyeball progress and likely reprocessing requirements for DR7. Non-survey Data Release: RPB has been pushing some 09A non-survey datasets through the system for preparation of catalogue release DBs. Astrogrid deployment & Data Analysis services: MAR and ETWS noted that the crash problem in the list-driven photometry service has now been fixed, but that the intermittent problems remain. ACTION: MSH to create a login for RSC on horus ACTION: ETWS to point RSC to an example command application that is failing in the list-driven photometry service Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================